OpenStudy (anonymous):

there are n teams in a league and the number of games played is given by the polynomial n^2-n=N. A womens softball league has 9 teams . what is the total numbers of games to be played

OpenStudy (saifoo.khan):

idk where, but i have done all these before....

OpenStudy (saifoo.khan):

(9^2)-9=N N = 81 - 9 N = 72